Suppose You have eight televisions. Three are defective. If teo are randomly selected, compute the probability that both televisions work. What is the probability that at least one of the two does not work?
1
Expert's answer
2012-10-04T08:31:31-0400
Since three of them are defective, 5 TV should be in good quality, the probability that none of the two television are defective would then be

5/8 * 4/7 = 5/14

and the probability that at least one TV is defective is

1 - 5/14 = 9/14.